### v1

# Surprisingly fragile: Assessing and addressing prompt instability in multimodal foundation models

Abstract
Multimodal foundation models (MFMs) such as OFASys show the potential to unlock analysis of complex data such as images, videos, and audio data via text prompts alone. However, their performance may suffer in the face of text input that differs even slightly from their training distribution, which is surprising considering the use of modality-specific data to “ground” the text input. This study demonstrates that prompt instability is a major concern for MFMs, leading to a consistent drop in performance across all modalities, but that instability can be mitigated with additional training with augmented data. We evaluate several methods for grounded prompt perturbation, where we generate perturbations and filter based on similarity to text and/or modality data. After re-training the models on the augmented data, we find improved accuracy and more stable performance on the perturbed test data regardless of perturbation condition, suggesting that the data augmentation strategy helps the models handle domain shifts more effectively. In error analysis, we find consistent patterns of performance improvement across domains, suggesting that retraining on prompt perturbations tends to help general reasoning capabilities in MFMs.
Introduction
While text-only large language models (LLMs) are powerful in their own right, multimodal foundation models offer even more transformative capabilities such as image captioning and audio signal processing (li2024). Generalist models such as Unified-IO are trained to process a range of modalities including image, video, audio, and depth maps, often using an LLM as a combined encoder for all input types  (lu2022). MFMs have the potential for major impact in a variety of downstream applications such as autonomous navigation (cui2024) and medical image analysis (moor2023), and many public-facing LLMs such as ChatGPT are increasingly incorporating multimodal modeling into the user experience to maximize their impact.
Despite their power, these models often show surprising fragility in the face of data that deviates slightly from the expected format: a model may respond very differently to a prompt that has a slight difference in phrasing (newman2021). We show an example from a popular video QA dataset (iVQA) in Table 1, in which a person fixes a car engine and the model is asked to respond to the prompt “What mechanism is the person working on?”. In response to prompt perturbations, the Unified-IO model with no additional training responds with a variety of incorrect responses, while the model that has been trained on prompt perturbations (“Joint similarity”; see Prompt augmentation) responds correctly regardless of phrasing.
Data Prompt Answer
[Uncaptioned image] What mechanism is the person working on? The person in the video seems to be working on an engine mechanism.
Perturbed prompt Model response (no training) Model response (trained on prompt perturbations)
what project is the person working on? home The person in the video is working with a car engine.
what technology is the person improving? oil The person in the video appears to be working on a car.
what task is the person in charge of? excavator The person in charge of the vehicle seems to be working on the engine.
Table 1: Unified-IO model responses to prompt perturbations on example data (iVQA video, first frame shown).
We investigate the relative stability of current state-of-the-art MFMs on several standard multimodal benchmark datasets and find the following:
  • Modern MFMs show a severe drop in performance in response to perturbed prompts on standard QA datasets (Testing on perturbed data).
  • Re-training generalist multimodal models on a wider variety of perturbed prompts, regardless of how the prompts are selected (Prompt augmentation), improves performance and reduces instability on other prompt perturbations, more than simply re-training on the original prompts from the task data (Training on perturbed data). The perturbation-trained models show consistent improvement when the prompts target modality-specific information, e.g. prompts in the image modality that reference material properties (Error analysis).
Our work provides an automated method to help align MFMs to human expectations, such as consistency in response to input variation (zhao2024).
Related work
Multimodal foundation models
Recent work has unified text-based foundation models with multimodal processing to yield multimodal foundation models that can respond to the combination of text and multimodal input (li2024). The majority of such models, including Unified-IO and OFASys, first encode the text and multimodal data to a shared representation space, then decode the combined representation into the desired modality, often text (e.g. image QA requires converting image + text to text). Some systems such as Next-GPT learn to perform any-to-any modality generation by training modality-specific decoders (wu2023), which works reasonably well for more common modalities such as images. For breadth, systems including ImageBind (girdhar2023) and OneLLM (han2023) use a universal encoder and learn from a wide variety of data including depth maps and point clouds, which may help the model’s internal LLM achieve better cross-modality understanding. To further improve performance, parameter-efficient methods such as prompt learning have helped to solve problems with missing data (lee2023) and cross-modality alignment (khattak2023). We identify a possible gap in the research space, i.e. model-agnostic methods to improve robustness, and test a variety of prompt perturbation methods to determine impact on trained model performance.
Model stability
Many LLMs struggle with limited stability: the response to a given input may vary widely depending on changes in input wording or semantics (newman2021), even as subtle as adding extra punctuation (formento2023). This not only presents a negative user experience to end users, requiring them to iterate on the same prompt (kim2024), but also open vulnerability to attacks that can lead to model output of harmful information (schulhoff2023). One common solution to this problem is model regularization, i.e. forcing the model to learn general patterns in the data, and methods for regularization include enforcing sparsity in model architecture (li2023), or more commonly augmenting training data with “noised” data points (chen2023). In the multimodal space, methods such as synonym replacement have shown to generate useful data points at inference time to reduce instability and improve overall performance (ge2023). We address the issue of stability through modality-grounded prompt perturbation and show that re-training on such prompts can help insulate models from performance drops on unseen prompts.
Methods
We first outline the methods used for sampling data-grounded prompts for further model training (Prompt augmentation), identify the models trained using the augmented data (Models), and explain the QA data used in the experiments (Evaluation protocol).
Prompt augmentation
Typical approaches to data augmentation include utilizing a variety of similar but distinct prompts to re-train the model, e.g. randomly changing the prompt instructions during training (lu2023). We test several simple methods for data augmentation to determine whether MFMs require modality-specific prompts or text-specific prompts to improve overall prompt stability.
For a given text prompt $t$, we generate a list of $N=10$ perturbations $\hat{T}$ with a general-purpose LLM, using instructions to maximize diversity and semantic similarity to the original prompt. [^1] Llama-2, accessed 1 Jan 2024: https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. We chose 10 as the number of perturbations to maximize prompt diversity without encouraging prompts that “drift” too far from the original prompt. This full set of perturbations $\hat{T}$ is used in evaluation to determine post-training performance of the models on the perturbed prompt data.
We then compute the latent representation $X$ of text ($X_{t}$) and multimodal data ($X_{m}$) using the ImageBind model (girdhar2023), an encoder trained to align latent representations of data from different modalities with contrastive learning, using text as the intermediate modality. Finally, we select $k=3$ prompts from the set $\hat{T}$ that have maximum cosine similarity between the prompt text and the latent representations of the text, choosing either $X_{t}$ for text-similar paraphrases (“text-similar sampling”) or $X_{m}$ for modality-similar paraphrases (“modality-similar sampling”). We also sample randomly from the prompt perturbations as a control condition (“random sampling”). We choose 3 as the number of prompts per category to balance semantic diversity with amount of runtime available for training.
In addition to these simpler conditions for prompt perturbation, we test a method to balance similarity of the perturbed prompt to the original data and diversity among the prompts. We sample 1 prompt perturbation $\hat{t}$ from $\hat{T}$ based on Eq. 1 similarity to original data (text and multimodal data).
[…]
Then, in order to increase the diversity of the jointly sampled prompts, we normalize the similarity relative to the cumulative similarity of the original prompt to the set of already-sampled prompts ($\hat{T}^{\prime}$) from prior iterations (Eq. 2). As each additional prompt perturbation $\hat{t}$ is sampled, it is added to $\hat{T}^{\prime}$.
[…]
This method (“Joint similarity”) yields a diverse set of input text that remains grounded with respect to the modality-specific data, thus decreasing the risk of re-training the model on augmented prompts that are irrelevant to the input context.
Prompt Data
What is the person holding? [Uncaptioned image]
Prompt perturbations
1. what is the person’s grasping tool?
2. what is the person’s hold?
3. what object is the person gripping?
4. what is the person’s holding device?
5. what is the person’s grasp?
6. what item is the person clutching?
7. what is the person’s grip?
8. what is the person’s clutch?
9. what object is the person grasping?
10. what is the person holding tightly?
Text similarity Modality similarity Joint-similarity
2. what is the person’s hold? 10. what is the person holding tightly? 4. what is the person’s holding device? 6. what item is the person clutching? 7. what is the person’s grip? 5. what is the person’s grasp? 8. what is the person’s clutch? 7. what is the person’s grip? 6. what item is the person clutching?
Table 2: Prompt sample methods from Prompt augmentation, with example data and prompt (video frame from iVQA).
We summarize the prompt sampling methods in Table 2, where we show an example data point with the corresponding sampled prompt perturbations.
Models
For our experiments, we tested image, video, and spoken audio data [^2] At time of writing, no open-source generalist models proved capable of handling general audio in addition to other modalities, only human speech audio., which are three related but complementary data modalities that a generalist multimodal model should reasonably be expected to handle. Despite the proliferation of literature related to MFMs, relatively few models are available for open-source inference, and even fewer for fine-tuning on existing model weights. We investigate the following models which have released open-source implementations including training code:
All the models encode multimodal data into the same latent representation space as text, provide text and multimodal data jointly to a universal processing unit, and generate output in an autoregressive fashion. We considered other models such as Macaw (lyu2023) and OneLLM (han2023) but did not adopt them due to the higher difficulty of implementation and seemingly poor performance outside of the image domain.
Evaluation protocol
We evaluate our proposed method for data sampling using the following QA datasets that have free-response questions:
  • VQA (antol2015): questions about MS COCO images
  • iVQA (yang2021): questions about how-to videos from YouTube
  • Spoken-SQuAD (lee18): questions about spoken Wikipedia articles (generated via text-to-speech)
We choose these tasks to approximate the kinds of tasks expected of MFMs, seen in the original work on OFASys and Unified-IO (lu2022; wang2022). We use a total of 1500 prompts and associated data (images, audio, video) for each dataset. Throughout evaluation (Results), we use the same 80%/20% train/test split of the data for fair comparison. Some of the models in this study had size limitations on the image and video data, which required extra data preprocessing: e.g., for images and videos, we pad and crop the data to 224 x 224 pixels; for videos, we sample 1-second frames for a total of 8 frames per video. We provide summary statistics about the data in Table TODO: summary statistics (description; question length; answer length; distribution of data size/length)
To match the open-response format of the QA tasks, we evaluate the models’ performance using typical text generation metrics including BLEU, ROUGE, and BERT-Score (zhang2020). In general, we find that the BERT-Score results have higher values than BLEU and ROUGE, due to the fuzzy nature of the metric (semantic similarity) as compared to the more strict word-overlap metrics that penalize slight changes in wording.
Results
Testing on perturbed data
We first assess the degree to which multimodal models are able to correctly respond to perturbed prompts generated via three separate models:
  1. 1.
    General-purpose LLM (see Prompt augmentation)
  2. 2.
    Decoder model fine-tuned on paraphrasing [^5] Model accessed 30 Aug 2023: https://huggingface.co/ramsrigouthamg/t5˙sentence˙paraphraser.
  3. 3.
    Back-translation model (between English and Russian) [^6] Models accessed 30 Aug 2023: https://huggingface.co/Helsinki-NLP/opus-mt-en-ru https://huggingface.co/Helsinki-NLP/opus-mt-ru-en.
We choose three different methods to assess a wide range of possible prompt changes that end users might make.
The results are shown in Table 3, in which the image modality has the most consistent drop in performance. We see consistent degradation of performance from the original questions to the perturbed prompts, up to a 50% drop in the case of Unified-IO and images (BLEU score of 0.687 for original prompts vs. 0.347 for LLM-paraphrase prompts). The video domain has slight degradation and low performance overall, due in part to the more complicated answers expected of the model (e.g., “The device being used in the video is a blender.”) as well as the more complex nature of videos in general. The overall low performance in the audio domain may be due to the inherently difficult nature of the QA task, which requires first converting speech to text and then performing logical reasoning over the text to answer the question. Surprisingly, both OFASys and Unified-IO reported high performance in audio and video tasks in the original work (e.g., up to 42.1% accuracy for OFASys on VQA for MSR-VTT data; bai2022 bai2022), which may indicate overfitting on a specific format of the task.
Based on this analysis, we conclude that prompt stability is a major concern for generalist multimodal models even in “strong” domains such as images, and a less substantial problem for less well-supported domains such as video.
OFASys Unified-IO
Modality Prompt type BLEU ROUGE BERT BLEU ROUGE BERT
Audio Original 0.0434 (0.0052) 0.1286 (0.0081) 0.8153 (0.0023) 0.0142 (0.0037) 0.0542 (0.0069) 0.8215 (0.0021)
Back-translation 0.0468 (0.0041) 0.1316 (0.0061) 0.8150 (0.00174) 0.0115 (0.0022) 0.0437 (0.0038) 0.8171 (0.0014)
Paraphrase 0.0431 (0.0029) 0.1288 (0.0047) 0.8151 (0.0014) 0.0113 (0.0015) 0.0484 (0.0032) 0.8143 (0.0010)
LLM-Paraphrase 0.0453 (0.0017) 0.1341 (0.0027) 0.8168 (0.0008) 0.0142 (0.0013) 0.0476 (0.0020) 0.8172 (0.0006)
Image Original 0.4647 (0.0271) 0.4739 (0.0270) 0.9646 (0.0036) 0.6878 (0.0248) 0.7012 (0.0245) 0.9808 (0.0026)
Back-translation 0.4053 (0.0233) 0.4145 (0.0233) 0.9648 (0.0031) 0.5551 (0.0233) 0.5664 (0.0232) 0.9743 (0.0026)
Paraphrase 0.3996 (0.0165) 0.4112 (0.0165) 0.9633 (0.0023) 0.4942 (0.0167) 0.5051 (0.0167) 0.9605 (0.0023)
LLM-Paraphrase 0.3832 (0.0089) 0.3928 (0.0089) 0.9636 (0.0012) 0.3470 (0.0087) 0.3563 (0.0087) 0.9272 (0.0016)
Video Original 0.0043 (0.0005) 0.0299 (0.0037) 0.8174 (0.0005) 0.0171 (0.0033) 0.0877 (0.0077) 0.8294 (0.0016)
Back-translation 0.0038 (0.0004) 0.0266 (0.0027) 0.8177 (0.0004) 0.0113 (0.0016) 0.0659 (0.0051) 0.8281 (0.0012)
Paraphrase 0.0037 (0.0003) 0.0259 (0.0021) 0.8173 (0.0003) 0.0193 (0.0023) 0.0658 (0.0046) 0.8257 (0.0010)
LLM-Paraphrase 0.0039 (0.0002) 0.0279 (0.0012) 0.8170 (0.0002) 0.0058 (0.0004) 0.0334 (0.0014) 0.8176 (0.0003)
Table 3: Performance of no-training model on original and perturbed prompts (standard error in parentheses).
Training on perturbed data
Training conditions
We address the performance degradation on perturbed data by re-training the MFMs using the prompts generated through data augmentation (Prompt augmentation), applied to the QA datasets. We follow the same training procedure across all models, using the same hyperparameters7773 training epochs, learning rate $55e-$ 5, batch size 2, Adam optimizer with default hyperparameters. and objective function of cross-entropy on the model output logit probabilities compared to the ground-truth data. For the original prompt condition, we only train the model on the original prompts from all QA datasets. [^8] We also tried a condition where we trained on the same number of original prompts as the sample prompts, i.e. repeat each original prompt $N=3$ times in the training data, and found substantially worse performance across models and tasks. The decoder takes 99.3M parameters for OFASys and 409M parameters for Unified-IO, which is sufficiently lightweight to train on a single A100 GPU.
For each prompt perturbation condition, we use 3 sampled prompts from that condition in place of the original prompt (e.g. top-3 prompts with modality data similarity for “Modality-similarity” condition). OFASys and Unified-IO have a similar architecture that uses an LLM as the decoder for the combined input from text and other modalities, and we only retrain this decoder in both models as we are primarily interested in improving stability in the generated output, i.e. directly from the decoder, rather than improving e.g. encoder representation stability.
To summarize, we evaluate the following training conditions for each model type:
  • No training
  • Train only on original prompt (“Original”)
  • Train on prompt perturbations (“Joint similarity”, “Modality-similarity”, “Text similarity”, “Random sample”; see Prompt augmentation).
Results
We first show the model results on the original prompts in Table 4. The models without training (“No training”) perform poorly on audio and video and reasonably well on image. The models trained on the original prompts perform slightly better than the models trained on the perturbed prompts for the image domain, worse for the audio domain, and roughly the same for the video domain. The models trained on random samples of prompt perturbations generally perform worse than the other perturbation-trained models (e.g., image domain for OFASys), which may be due to noisy prompts that have little semantic relation with the meaning of the original prompt. In general, the results don’t show evidence of catastrophic forgetting, where the models re-trained on prompt perturbations would lose their ability to understand the original prompts due to overfitting on the perturbations.
OFASys Unified-IO
Modality Training condition BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0434 (0.0052) 0.1286 (0.0081) 0.8153 (0.0023) 0.0142 (0.0037) 0.0542 (0.0069) 0.8215 (0.0021)
Original prompts 0.1671 (0.0193) 0.2061 (0.0209) 0.8882 (0.0043) 0.0881 (0.0144) 0.1188 (0.0157) 0.8629 (0.0037)
Joint similarity 0.3749 (0.0266) 0.4034 (0.0267) 0.9131 (0.0049) 0.2016 (0.0218) 0.2248 (0.0222) 0.8810 (0.0043)
Modality-similarity 0.3708 (0.0266) 0.3982 (0.0268) 0.9125 (0.0048) 0.2205 (0.0227) 0.2440 (0.0230) 0.8883 (0.0044)
Text-similarity 0.3910 (0.0269) 0.4169 (0.0269) 0.9166 (0.0049) 0.2303 (0.0231) 0.2526 (0.0234) 0.8857 (0.0045)
Random sample 0.3445 (0.0260) 0.3743 (0.0263) 0.9095 (0.0048) 0.2089 (0.0223) 0.2305 (0.0227) 0.8813 (0.0044)
Image No training 0.4647 (0.0271) 0.4739 (0.0270) 0.9646 (0.0036) 0.6878 (0.0248) 0.7012 (0.0245) 0.9808 (0.0026)
Original prompts 0.7102 (0.0243) 0.7188 (0.0240) 0.9812 (0.0027) 0.7400 (0.0237) 0.7507 (0.0233) 0.9829 (0.0025)
Joint similarity 0.6650 (0.0255) 0.6733 (0.0253) 0.9791 (0.0029) 0.7186 (0.0241) 0.7306 (0.0237) 0.9836 (0.0024)
Modality-similarity 0.6494 (0.0257) 0.6589 (0.0255) 0.9767 (0.0030) 0.7208 (0.0242) 0.7334 (0.0239) 0.9834 (0.0025)
Text-similarity 0.6463 (0.0258) 0.6572 (0.0256) 0.9738 (0.0032) 0.7401 (0.0237) 0.7510 (0.0234) 0.9845 (0.0025)
Random sample 0.6189 (0.0260) 0.6294 (0.0259) 0.9734 (0.0032) 0.7130 (0.0243) 0.7225 (0.0240) 0.9835 (0.0025)
Video No training 0.0043 (0.0005) 0.0300 (0.0037) 0.8174 (0.0005) 0.0171 (0.0033) 0.0877 (0.0077) 0.8294 (0.0016)
Original prompts 0.4524 (0.0140) 0.6499 (0.0116) 0.9509 (0.0017) 0.4927 (0.0160) 0.6703 (0.0130) 0.9539 (0.0018)
Joint similarity 0.4472 (0.0155) 0.6285 (0.0129) 0.9485 (0.0019) 0.4597 (0.0176) 0.6302 (0.0146) 0.9467 (0.0021)
Modality-similarity 0.4483 (0.0156) 0.6286 (0.0128) 0.9483 (0.0019) 0.4507 (0.0170) 0.6231 (0.0142) 0.9461 (0.0021)
Text-similarity 0.4683 (0.0159) 0.6474 (0.0127) 0.9513 (0.0018) 0.4914 (0.0183) 0.6537 (0.0147) 0.9499 (0.0021)
Random sample 0.4387 (0.0152) 0.6232 (0.0125) 0.9478 (0.0018) 0.4503 (0.0177) 0.6250 (0.0147) 0.9464 (0.0021)
Table 4: Model accuracy on original prompts (standard error in parentheses).
Next, we evaluate model performance on all prompt perturbations ($\hat{T}$) generated for data augmentation (Prompt augmentation), to determine whether one method of prompt perturbation yields better generalization. The results are shown in Table 5. Across all modalities, the baseline model and the model trained on original prompts show substantial decrease in performance on prompt perturbations versus the original prompts. In contrast, the models trained on the prompt perturbations show consistently high performance on the prompt perturbations, with a slight decrease in the video modality and an increase in the audio modality. Surprisingly, the models show consistent performance regardless of the prompt sampling method used, which implies that simply exposing the model to a wider diversity of prompts always helps in the perturbation test setting, even if the exact form of the prompts varies. In the video domain, the regular prompt-trained model has substantially worse performance than the perturbation-trained models, possibly due to a more limited semantic space covered by the original prompts in the video domain as compared to the prompt perturbations (e.g. covering different verbs to describe the same motion or actions taken in the videos).
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0453 (0.0017) 0.1341 (0.0027) 0.8168 (0.0008) 0.0142 (0.0013) 0.0476 (0.0020) 0.8172 (0.0006)
Original prompts 0.1186 (0.0053) 0.1565 (0.0060) 0.8741 (0.0014) 0.0427 (0.0030) 0.0715 (0.0036) 0.8531 (0.0011)
Joint similarity 0.4367 (0.0088) 0.4701 (0.0088) 0.9208 (0.0016) 0.2128 (0.0074) 0.2316 (0.0074) 0.8820 (0.0014)
Modality-similarity 0.4390 (0.0089) 0.4714 (0.0088) 0.9206 (0.0016) 0.2230 (0.0075) 0.2415 (0.0076) 0.8832 (0.0015)
Text-similarity 0.4542 (0.0089) 0.4860 (0.0088) 0.9223 (0.0016) 0.2317 (0.0076) 0.2507 (0.0077) 0.8848 (0.0015)
Random sample 0.4090 (0.0087) 0.4453 (0.0087) 0.9172 (0.0015) 0.2311 (0.0076) 0.2494 (0.0077) 0.8835 (0.0015)
Image No training 0.3832 (0.0089) 0.3928 (0.0089) 0.9636 (0.0012) 0.3470 (0.0087) 0.3563 (0.0087) 0.9272 (0.0016)
Original prompts 0.4572 (0.0091) 0.4647 (0.0091) 0.9668 (0.0012) 0.3754 (0.0088) 0.3851 (0.0088) 0.9386 (0.0019)
Joint similarity 0.6963 (0.0084) 0.7002 (0.0084) 0.9796 (0.0010) 0.6371 (0.0088) 0.6415 (0.0088) 0.9791 (0.0010)
Modality-similarity 0.6912 (0.0085) 0.6952 (0.0084) 0.9793 (0.0010) 0.6470 (0.0088) 0.6512 (0.0087) 0.9800 (0.0009)
Text-similarity 0.6885 (0.0085) 0.6941 (0.0084) 0.9795 (0.0010) 0.6226 (0.0089) 0.6280 (0.0088) 0.9771 (0.00100)
Random sample 0.6909 (0.0085) 0.6961 (0.0084) 0.9791 (0.0010) 0.6495 (0.0088) 0.6532 (0.0087) 0.9801 (0.0009)
Video No training 0.0039 (0.0002) 0.0279 (0.0012) 0.8170 (0.0002) 0.0058 (0.0004) 0.0334 (0.0014) 0.8176 (0.0003)
Original prompts 0.0238 (0.0013) 0.0730 (0.0025) 0.8262 (0.0005) 0.1390 (0.0034) 0.2720 (0.0049) 0.8526 (0.0028)
Joint similarity 0.3857 (0.0047) 0.5760 (0.0042) 0.9414 (0.0006) 0.4301 (0.0063) 0.5878 (0.0052) 0.9417 (0.0008)
Modality-similarity 0.3824 (0.0047) 0.5752 (0.0041) 0.9412 (0.0006) 0.4394 (0.0065) 0.5947 (0.0053) 0.9422 (0.0008)
Text-similarity 0.3854 (0.0047) 0.5765 (0.0041) 0.9416 (0.0006) 0.4378 (0.0064) 0.5937 (0.0053) 0.9424 (0.0008)
Random sample 0.3825 (0.0047) 0.5744 (0.0041) 0.9413 (0.0006) 0.4401 (0.0064) 0.5936 (0.0053) 0.9426 (0.0008)
Table 5: Model accuracy on prompt perturbations (standard error in parentheses).
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 2.0588 1.0597 0.0506 5.0223 2.2407 0.0401
Original prompts 2.3907 2.0382 0.0838 3.8096 2.7091 0.0691
Joint similarity 1.0825 1.0023 0.0908 1.860 1.7236 0.0874
Modality-similarity 1.0811 1.0024 0.0905 1.8092 1.6807 0.0883
Text-similarity 1.0484 0.9740 0.0901 1.7637 1.6400 0.0882
Random sample 1.1384 1.0469 0.0900 1.7672 1.6517 0.0927
Image No training 1.2503 1.2221 0.0689 1.3487 1.3197 0.0930
Original prompts 1.0750 1.0576 0.0659 1.2685 1.2372 0.1103
Joint similarity 0.6543 0.6473 0.0545 0.7499 0.7413 0.0533
Modality-similarity 0.6630 0.6558 0.0549 0.7333 0.7255 0.0521
Text-similarity 0.6657 0.6551 0.0550 0.7708 0.7607 0.0555
Random sample 0.6632 0.6536 0.0552 0.7300 0.7231 0.0519
Video No training 2.2890 2.2632 0.0103 3.7975 2.3034 0.0183
Original prompts 3.0050 1.8506 0.0354 1.3150 0.9801 0.1785
Joint similarity 0.6646 0.3932 0.0348 0.7952 0.4805 0.0437
Modality-similarity 0.6595 0.3842 0.0345 0.7952 0.4815 0.0446
Text-similarity 0.6623 0.3890 0.0347 0.7906 0.4794 0.0441
Random sample 0.6629 0.3909 0.0348 0.7853 0.4817 0.0441
Table 6: Model coefficient of variation, on prompt perturbations (lower is better).
To assess the consistency of the performance differences over the prompt perturbations test data, we compute the coefficient of variation $CV$ ($\frac{\sigma^{2}}{\mu}$) over all the prompt perturbations, i.e. compute the standard deviation in performance score over all prompts for a given dataset/model and normalize by the mean. The normalized value represents the amount of variation in score relative to the expected score, e.g. a standard deviation of 5 in BLEU score for a high-scoring model should be better than the same standard deviation for a low-scoring model.
Shown in Table 6, the re-trained models achieve lower $CV$ on the prompt perturbations. This shows that training with any of the prompt perturbation strategies yields more stable performance as compared to training only on the original prompts. The variation is slightly higher in the audio domain, which aligns with the generally lower performance on all forms of prompt perturbations in audio (Table 3), such that even a small absolute error is relatively large in comparison to the overall lower performance.
OFASys Unified-IO
Modality Cluster theme Data description Prompt Perturb prompt True answer Joint similarity Original prompts Joint similarity Original prompts
Audio Religion Spoken: “In the United States, the UMC ranks as the largest mainline Protestant denomination, the largest Protestant church after the Southern Baptist Convention, and the third largest Christian denomination.” In the United States the UMC ranks as the largest what? in the united states, the umc ranks as the largest _______ of christian organizations. mainline protestant denomination largest protestant christian mainline protestant denomination national
Image Home activities Well-lit kitchen with a table, preparation station, counters, cabinets, oven, and sink. What is the table made of? what does the table primarily consist of in terms of substance? marble marble glass marble table
Video Personal grooming Person with long hair sitting in a room, brushing their hair. How was the girl combing her hair? how did she groom her hair? The girl in the video was combing her hair with a brush. The woman in the video combed her hair with a brush. dry The girl in the video was brushing her hair with a brush. she brushed her hair
Table 7: Example data from target modalities, clustered by semantic representation. Cluster themes determined by manual inspection of prompts. Models organized by architecture and by training condition: “Original prompts” indicates models trained only original prompts, “Joint similarity” indicates models trained on prompt perturbations filtered for similarity to modality data and text data.
BLEU (mean)
Modality Cluster Example prompts Perturbation- training (all models) Original prompts Ratio
Audio 0 (religion) What do progressive moderates of Islam seek to separate? In the United States the UMC ranks as the largest what? What did the new mass allow as regards ceremony? 0.4755 0.0347 13.70
2 (space) What happened to the Apollo 13 crew to force them to use the LM as a ”life boat” as it had been designed? What is the name for O3 most often used? What did NASA name the series seeking up to 30 more flights to Earth’s orbit? 0.5583 0.0467 11.96
4 (law, history) What have the Treaties sought to enable since its foundation? Which courts have a duty to interpret domestic law as far as possible? How can you protest against the government in an individual way? 0.5300 0.0758 6.99
Image 3 (street views) What is on the sidewalk beside the woman? How many signs are on the pole? What is on the ladder rungs? 0.9050 0.3713 2.44
4 (people eating together) Where are these people eating? What color is the person’s shirt? Does this look more like a party or a funeral? 0.7344 0.3912 1.88
7 (Home activities) What color are the walls? What kind of shop is this? How many bags are black? 0.7049 0.3851 1.83
Video 2 (personal objects) What instrument is the person playing? What is the person wearing on her neck? What type chair is the person sitting on? 0.4131 0.0126 32.79
3 (hand movement) What is placed next to the bowl? What is she holding? What does the subject slide into the machine? 0.5262 0.0177 29.73
0 (personal grooming) What tool did she use? How was the girl combing her hair? What promenant fashion accessory is the lady wearing? 0.4195 0.0156 26.89
Table 8: Per-modality clusters and model scores, for perturbation-trained and prompt-trained OFASys models (other models omitted for space), sorted by ratio of BLEU score on prompt perturbations (mean for all perturbation-trained models, divided by prompt-trained model).
Error analysis
We perform a systematic error analysis to determine possible reasons for the performance improvement on the perturbed prompts (Table 5), comparing the perturbation-trained models to the prompt-trained models. We first embed all the modality data using the ImageBind model (Prompt augmentation), reduce to $D=3$ dimensions with PCA, and then cluster data from each modality via cosine similarity of the embeddings. The clustering was implemented with a hierarchical DBSCAN with default hyperparameters (minimum cluster size 5, automatic algorithm selection) which identified $K={15,11,12}$ clusters for audio, image, and video modalities respectively. We identify the dominant concept in each cluster based on the associated prompts, and identify clusters with an especially high improvement from the perturbation-trained models as compared to the prompt-trained models. The high-improvement clusters are shown in Table 7, e.g., the image cluster related to street views shows a roughly 250% improvement from prompt-trained to perturbation-trained (0.3713 to 0.9050).
For audio, the example cluster represents spoken explanations of religious organizations, such as the UMC (United Methodist Church). This topic of data likely requires more complicated and contextualized answers to questions, which the perturbation-trained models seem able to provide (e.g. OFASys response, “largest protestant”) while the models trained only on original prompts produce more short and ambiguous responses (OFASys, “national”). For image, the example cluster contains scenes of home activities, including a typical kitchen scene. These data include prompts that ask about basic object properties such as quantity and material type, and the perturbation-trained models appear more able to understand the intent of the prompts as they relate to properties such as “substance” (e.g., Unified-IO, “marble”). For video, the example cluster includes clips of people facing the camera, mainly performing tasks with tools such as applying makeup and combing their hair. The prompts for this data ask for further descriptions of the actions shown, and the prompt-trained models have difficulty understanding the full context of differently-phrased actions (e.g., OFASys, “dry”) as compared to the perturbation-trained models are able to provide a full description of the actions (OFASys, “The woman in the video combed her hair with a brush.”).
We show the top-3 clusters per modality in Table 8 to show more example prompts and aggregate model scores. The improvement across highly diverse topics and contexts shows that the perturbation-trained models’ improvements are not spurious (e.g., only applying to a specific type of QA task) but reflect an improvement in general reasoning ability.
Conclusion
This study showed that multimodal foundation models (MFMs) show a surprising degree of instability in response to prompt perturbations, and that the instability can be mitigated through fine-tuning on additional prompt perturbations, regardless of whether the perturbations are close to the original prompt or modality data. The finding about re-training on prompt perturbations is surprising but suggests that MFM developers should focus more on diversity of prompts rather than prompt similarity to the original data, which appears to be less relevant for models during training. Lastly, the error analysis shows systematic improvements in specific clusters of data in different modalities, e.g. in videos that depict personal grooming, and this suggests that the additional training helps reinforce modality-specific properties to the model that leads to greater robustness.
Future work may investigate more complicated paradigms for increasing model robustness, including direct preference optimization to determine the relative value of the different perturbations for model training (rafailov2024; zhao2024). In this work we only considered perturbations of the text data, but future work should test model robustness in response to perturbations of the modality data, e.g., adding additional noise to images (lee2020). Lastly, future work should consider more structured/explainable methods for determining similarity between modality data and text: instead of ImageBind, one may use the texts output from a well-trained image caption model to compare against a prompt perturbation to determine if the perturbation has “drifted” too far from the image content.
Acknowledgments
This project was funded by NA-22…TODO: funding details
Limitations
The prompt perturbations in this work are meant to emulate real-world variation in prompts due to differences in user input. However, this is a strong assumption that may not hold in practice, and we cannot guarantee that the perturbations generated in this study are a perfect match for user input. We hope that future work can investigate this mismatch more closely and determine the best proxy for variation in real-world prompts. Ideally, this work would provide a lower bound for the possible diversity in prompt space due to differences among system users, considering that human behavior with respect to prompting LLMs can be somewhat unpredictable (kim2024).
Appendix A Results by prompt perturbation type
We further analyze the aggregate prompt perturbation results (Table 5) by sub-dividing them by prompt perturbation type, i.e. joint similarity, modality-similarity, and text-similarity. The results per-type are shown in separate tables: text-similarity prompt results in Table 9, modality-similarity prompt results in Table 10, and joint similarity prompt results Table 11.
The prompt-perturbation models have higher accuracy for their respective perturbation type, e.g. for the text-similarity prompts the model with highest performance was the model trained on text-similarity prompts. This suggests that the prompt perturbation filtering processes tend to generate self-similar prompts, which leads to more consistent performance from the models that are trained on those filtered prompts when applied to those self-similar prompts. The self-similarity is apparent even in the example prompts in LABEL:fig:prompt_sample_method, where the text-similarity prompt perturbations focus on the act of “holding” rather than the object being held or the person doing the holding.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0450 0.1337 0.8169 0.0145 0.0469 0.8170
Original prompts 0.1279 0.1630 0.8774 0.0449 0.0709 0.8521
Joint similarity 0.4486 0.4797 0.9225 0.2202 0.2399 0.8842
Modality-similarity 0.4518 0.4853 0.9233 0.2350 0.2519 0.8859
Text-similarity 0.4941 0.5239 0.9274 0.3158 0.3320 0.8976
Random sample 0.4138 0.4492 0.9186 0.2432 0.2604 0.8850
Image No training 0.3947 0.4045 0.9644 0.3827 0.3920 0.9308
Original prompts 0.5026 0.5107 0.9699 0.4083 0.4193 0.9423
Joint similarity 0.7308 0.7339 0.9807 0.6673 0.6731 0.9824
Modality-similarity 0.7253 0.7287 0.9802 0.6848 0.6891 0.9837
Text-similarity 0.7767 0.7809 0.9840 0.7383 0.7432 0.9856
Random sample 0.7232 0.7280 0.9806 0.6870 0.6921 0.9835
Video No training 0.0045 0.0314 0.8170 0.0062 0.0364 0.8177
Original prompts 0.0230 0.0745 0.8250 0.1588 0.2967 0.8562
Joint similarity 0.4182 0.6011 0.9451 0.4675 0.6224 0.9463
Modality-similarity 0.4177 0.6037 0.9452 0.4892 0.6326 0.9478
Text-similarity 0.4402 0.6204 0.9478 0.6097 0.7179 0.9599
Random sample 0.4167 0.6009 0.9455 0.4681 0.6159 0.9461
Table 9: Model accuracy on text-similarity perturbations.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0446 0.1335 0.8167 0.0126 0.0456 0.8175
Original prompts 0.1376 0.1787 0.8750 0.0464 0.0768 0.8529
Joint similarity 0.4300 0.4646 0.9190 0.2178 0.2404 0.8830
Modality-similarity 0.4670 0.4984 0.9241 0.2825 0.2978 0.8923
Text-similarity 0.4523 0.4867 0.9213 0.2366 0.2571 0.8854
Random sample 0.4123 0.4499 0.9175 0.2201 0.2402 0.8811
Image No training 0.3753 0.3854 0.9642 0.3439 0.3538 0.9266
Original prompts 0.4602 0.4696 0.9653 0.3770 0.3868 0.9442
Joint similarity 0.6971 0.7013 0.9795 0.6330 0.6374 0.9786
Modality-similarity 0.7394 0.7431 0.9826 0.7195 0.7222 0.9849
Text-similarity 0.6913 0.6980 0.9805 0.6322 0.6374 0.9772
Random sample 0.6902 0.6959 0.9806 0.6391 0.6439 0.9802
Video No training 0.0041 0.0292 0.8168 0.0068 0.0360 0.8172
Original prompts 0.0219 0.0694 0.8246 0.1516 0.2843 0.8610
Joint similarity 0.3980 0.5859 0.9429 0.4501 0.6041 0.9439
Modality-similarity 0.4132 0.6007 0.9449 0.5710 0.6863 0.9558
Text-similarity 0.3995 0.5886 0.9433 0.4706 0.6188 0.9462
Random sample 0.3958 0.5858 0.9430 0.4403 0.5945 0.9430
Table 10: Model accuracy on modality-similarity perturbations.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0448 0.1336 0.8168 0.0154 0.0496 0.8169
Original prompts 0.1232 0.1585 0.8755 0.0439 0.0725 0.8533
Joint similarity 0.4625 0.4949 0.9247 0.2597 0.2739 0.8883
Modality-similarity 0.4379 0.4678 0.9210 0.2268 0.2438 0.8824
Text-similarity 0.4474 0.4802 0.9211 0.2353 0.2543 0.8843
Random sample 0.3959 0.4292 0.9154 0.2292 0.2461 0.8823
Image No training 0.3936 0.4047 0.9636 0.3497 0.3570 0.9289
Original prompts 0.4556 0.4630 0.9670 0.3764 0.3868 0.9362
Joint similarity 0.7399 0.7447 0.9805 0.7223 0.7263 0.9837
Modality-similarity 0.7027 0.7075 0.9798 0.6711 0.6757 0.9795
Text-similarity 0.6871 0.6937 0.9782 0.6396 0.6455 0.9780
Random sample 0.6908 0.6961 0.9788 0.6649 0.6684 0.9805
Video No training 0.0041 0.0287 0.8169 0.0067 0.0352 0.8171
Original prompts 0.0225 0.0737 0.8259 0.1496 0.2871 0.8628
Joint similarity 0.4106 0.5958 0.9439 0.5379 0.6669 0.9526
Modality-similarity 0.3941 0.5866 0.9424 0.4536 0.6043 0.9440
Text-similarity 0.3947 0.5836 0.9429 0.4587 0.6099 0.9447
Random sample 0.3932 0.5805 0.9424 0.4499 0.6003 0.9438
Table 11: Model accuracy on joint similarity perturbations.